OpenStudy (spitfire77):

suppose x is a random variable with mean 20 and standard deviation 5. also suppose that Y is a random variable with mean 40 and standard deviation 10. find the variance and standard deviation of the random variable z A) Z = 2 + 10x B) Z = X + Y

OpenStudy (spitfire77):

I have no idea how to start this

OpenStudy (anonymous):

Okay, let's start with notations. \[\Large \mathbb{E}\] means "mean" \[\Large \text{Var}\] means Variance. It would help you to know that \[\Large \mathbb{E} [X + Y]=\mathbb{E}[X]+\mathbb{E}[Y]\] and that \[\Large \mathbb E [kX]=k\mathbb E [X]\]

OpenStudy (anonymous):

Also, for Variances, we have \[\Large \text{Var}[X+Y] = \text{Var}[X] + \text{Var}[Y]\] and \[\Large \text{Var}[kX] = k^{\color{red}2}\text{Var}[X]\] Note the exponent.

OpenStudy (spitfire77):

what will the k equal in this case

OpenStudy (anonymous):

k is any constant. Now, here's the deal, I'll demonstrate the first item to you, then you'll demonstrate the second item to me. Sound fair?

OpenStudy (spitfire77):

yes

OpenStudy (anonymous):

Okay, let's begin. \[\Large Z = 2 + 10X\]We need its mean. So... \[\Large \mathbb E [Z] = \mathbb E [2 + 10X]\] By the first property of means, we get \[\Large \mathbb E[Z] = \mathbb E[2] + \mathbb E [10X]\]

OpenStudy (anonymous):

Now the mean of a constant is itself, and the variance of a constant is zero (rightly so, constants don't CHANGE, hence the name) So... \[\Large \mathbb E [Z] = \color{red}2 + \color{blue}{10}\mathbb E[X]\] by the second property of means ( the one that says E[kX] = k E[X] And we're almost done here, since we know that the mean of X is 20. We can just replace E[X] with 20. \[\Large \mathbb E[Z] = 2 + 10(\color{green}{20})\] \[\Large \mathbb E[Z] = 2 + 200 = 202\] And that's your mean.

OpenStudy (anonymous):

Questions?

OpenStudy (spitfire77):

so will the SD affect anything? or is it as simple as that

OpenStudy (anonymous):

Not yet, we have only ever tried to find the mean. Now we find the standard deviation. Let's do that by getting the variance, first. \[\Large Z = 2 + 10X\] \[\Large \text{Var}[Z] = \text{Var}[2+10X]\] By the first property of Variances, we get \[\Large \text{Var}[Z] = \text{Var}[2] + \text{Var}[10X]\]

OpenStudy (anonymous):

Now, the Variance of 2... a constant, since 2 doesn't VARY at all, is zero. So we are left with \[\Large \text{Var}[Z] = \text{Var}[10X]\] By the second property of Variances, 10 can move out, but becomes squared in the process: \[\Large \text{Var}[Z] = 10^{\color{red}2}\text{Var}[X] = 100\text{Var}[X]\]

OpenStudy (anonymous):

Now, since the standard deviation of X is 5, it follows that its variance is 25, since standard deviation is the square root of the variance. \[\Large \text{Var}[Z] = 100(25)\]

OpenStudy (spitfire77):

wasnt the sd 5? so the varience would be 25

OpenStudy (spitfire77):

ahh

OpenStudy (anonymous):

That's right :)

OpenStudy (anonymous):

Now, all that's left is to simplify \[\Large \text{Var}[Z] = 2500\] The Variance of Z is 2500, so to get the standard deviation, simply get the square root of the variance, and you'll get the SD of Z is 50

OpenStudy (spitfire77):

So Z has a var of 2500 and a sd of 50

OpenStudy (spitfire77):

ah beat me to it

OpenStudy (anonymous):

yup. So much for the first item. Now do the second one :)

OpenStudy (anonymous):

LOL We agreed that *I* was to do the first item and you do the second :D

OpenStudy (spitfire77):

so for B, it is Z = (20+40) Z has a mean of 60

OpenStudy (spitfire77):

then Z = 60(5) + 60(10) or is it 60(5+10)

OpenStudy (anonymous):

Correct. The SD?

OpenStudy (anonymous):

No... it's not that simple :P Get their variances :D

OpenStudy (spitfire77):

ok, give me a min

OpenStudy (spitfire77):

so the Var is 900 (Z)

OpenStudy (spitfire77):

30?

OpenStudy (spitfire77):

now i have another one, Z=10x-2 is it the same answer as A)?

OpenStudy (spitfire77):

also is the first part why didnt we multiply the mean (202) by the variance? like in the second question

OpenStudy (zarkon):

Are X and Y are independent?